OpenStudy (anonymous):

find the equation. through (3,8); parallel to f(x)=4x-2

OpenStudy (anonymous):

still (3,8) in point slope form with the slope of 4. y=8+4(x-3) y=8+4x-12 y=4x-4

OpenStudy (anonymous):

stick*

OpenStudy (anonymous):

wat is point slope form

OpenStudy (anonymous):

y=y1+m(x-x1) where (x1,y1) is a point and m is your slope